OpenStudy (anonymous):

Given vectors a and b, ware a = 5, b = 8 and angel between them is 60 degrees. calculate a . b (a DOT b)

OpenStudy (anonymous):

Here is the acutal question...

OpenStudy (turingtest):

the dot product of two vectors is given by the magnitudes times the cosine of the angle between them\[\vec A\cdot\vec B=|A||B|\cos\theta\]you are given all the info, so just plug the numbers into the formula

OpenStudy (turingtest):

*the product of the magnitudes times the cosine of the angle between them I mean

OpenStudy (anonymous):

ok so i get 20 .. and the next question says to do it b dot a... so do i just switch the a and b in the formula?

OpenStudy (anonymous):

becaosue i go the same answer and that dosent feel right

OpenStudy (anonymous):

a . b = 20 and b . a = 20

OpenStudy (turingtest):

yes, dot products are commutative\[\vec A\cdot B=\vec B\cdot\vec A\]so they should be the same

OpenStudy (anonymous):

and that would be the same even if i change the angle?

OpenStudy (turingtest):

if you change the angle, theta in your formula changes, so that will change the value of the dot product

OpenStudy (turingtest):

but it would still be commutative if that's what you mean

OpenStudy (anonymous):

but if the angle is same between b . a and a . b then the shuld be the same right?

OpenStudy (turingtest):

yes, the dot product is always commutative

OpenStudy (anonymous):

even if the angle is 90 both answers shuld be 90.. so =

OpenStudy (turingtest):

what is the cosine of 90 degrees?

OpenStudy (anonymous):

0 so it would cause everything to be 0

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

cool, you are awesome! you just clarified somuch for me... thx.. I wish i had someone like you for my entire assignment...

OpenStudy (turingtest):

Thanks, and you're welcome! I'm not hard to find here, and there are plenty of good tutors on this site welcome to OS, hope to see you around

OpenStudy (anonymous):

I will definately stay around... I was looking through many sites for help and got none and i saw an ad for OS and i came here and registed.

OpenStudy (anonymous):

for this one (see attachment) when i do a . a and b .b i get totally different results so what can i conclude about U . U ? - I dont even understand what they are asking.